OpenStudy (anonymous):

convert binery number to decimal 1001001

OpenStudy (anonymous):

after quick calcs 1001001=73

OpenStudy (anonymous):

Basically every number in a binary string represents a mathematical equation. For reach digit in the binary sequence, you take the value (0 or 1, let's call it n), and working from right to left, where i represents the index of the digit in the sequence you're looking at (starting at i = 0) \[n \times 2^{i}\] So, for your string, starting on the right we have 1, which is: \[1\times2^{0}=1\] Then 0: \[0\times2^{1}=0\] Then 0: \[0\times2^{2}=0\] Then 1: \[1\times2^{3}=8\] Then 0: \[0\times2^{4}=0\] Then 0: \[0\times2^{5}=0\] Then 1: \[1\times2^{6}=64\] Adding up all the resultant numbers gives us 1 + 8 + 64 = 73.

OpenStudy (anonymous):

You can get more information at http://www.wikihow.com/Convert-from-Binary-to-Decimal

OpenStudy (anonymous):

Rhys that is very confusing for any newbies to this. Here is a much simpler way of converting binary to decimal. 1, 2, 4, 8, 16, 32, 64 Notice the way I started with 1 and that doubled to a 2. The 2 then doubled to a 4. The 4 to a 8, 8 to 16 etc. Put them in reverse order and put your binary underneath 64, 32, 16, 8, 4, 2, 1 1, 0, 0, 1, 0, 0, 1 Therefore 64 + 8 + 1 = 73 Hope this helps.

OpenStudy (anonymous):

-55 is the answer if you're working with a signed binary number, 73 if you're unsigned.